Simon Companys year-end balance sheets follow. At December 31 Current Yr 1 Yr Ago 2 Yrs Ago $ 31,900 $ 34,700 $ 37,200 88,60For Year Ended December 31 Sales Cost of goods sold Other operating expenses Interest expense Income tax expense Total costsComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 2a Required 3 Compute the rComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 2a Required 3 Compute the pComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 2a Required 3 Assuming SimoComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 2a Required 3 Compute the d
Answer #1
1 Return on common stockholder's equity=Net income/Average common stockholder's equity
Average common stockholder's equity=(Beginning common stockholder's equity+Ending common stockholder's equity)/2
Common stockholder's equity=Common stock+Retained earnings
Current year 1 year ago
Net income a 37050 30190
Beginning common stock b 162500 162500
Beginning retained earnings c 94941 38622
Beginning common stockholder's equity d=b+c 257441 201122
Ending common stock e 162500 162500
Ending retained earnings f 156643 94941
Ending common stockholder's equity g=e+f 319143 257441
Average common stockholder's equity h=(d+g)/2 288292 229281.5
Return on common stockholder's equity i=a/h 12.85% 13.17%
2 Price earnings ratio=Common stock market price/Earnings per share
Earnings per share=Net income/Number of common shares outstanding
Number of common shares outstanding=Common stock in $/Par value=162500/10=16250
Current year 1 year ago
Common stock market price a 29 27
Net income b 37050 30190
Number of common shares outstanding c 16250 16250
Earnings per share d=b/c 2.28 1.86
Price earnings ratio e=a/d 12.72 14.53
2a. A higher PE ratio shows that investors are willing to pay a higher share price because of the growth expectation in the future.
Here, Simon has higher PE ratio than the competitor
Hence,investors have greater expectation about future performance of Simon.
3 Dividend yield=Annual cash dividend per share/Common stock market price
Current year 1 year ago
Annual cash dividend per share a 0.32 0.16
Common stock market price b 29 27
Dividend yield a/b 1.10% 0.59%